# Symbolic Music Generation with Non-Differentiable Rule Guided Diffusion

### v2

@@ 1 Introduction
We are interested in developing methods for controllable symbolic music generation. There has been rapid progress in the development of modern generative models for symbolic music (Huang et al., 2018; Huang & Yang, 2020; Hsiao et al., 2021; Min et al., 2023). To facilitate interaction between human composers and these models, it is crucial for these models to adhere to specific musical rules, such as chord progression, during the composition process. Moreover, these rules can be quite nuanced (e.g., the difference between a major chord and minor chord is very small).
@@ 2 Related Works
The conceptualization of stochastic optimal control in diffusion models has spurred theoretical advancements and practical applications. Zhang & Chen (2021) employed celebrated path integral theory to transform a simple Ornstein–Uhlenbeck process to a novel process whose target distribution matches given marginal distribution. Further extending this framework,  Berner et al. (2022); Vargas et al. (2023) established a novel link between stochastic optimal control problems and generative models, interconnected through stochastic differential equations.
@@ 3 Background
Classifier-free guidance (Ho & Salimans, 2022) avoids training classifiers by jointly training conditional and unconditional diffusion models, and combining their score estimates during sampling. The mixed score function becomes $(1+\omega)\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})-\omega\log p(\mathbf{x}_{t})$, where $\omega$ is the guidance strength. Despite easy implementation, it is expensive to extend classifier-free guidance to unknown or composite labels, because it requires re-training the diffusion model.
@@ 4.1 Rule Guidance Problem
A central challenge that we tackle is that many musical rules are non-differentiable, which makes sampling from Eq. 4 difficult. For instance, let $\mathbf{x}=[x_{1},x_{2},...,x_{n}]\in[0,1]^{n}$ be a vector where each $x_{i}$ represents the volume of a note, so that the note density is computed as $\operatorname{ND}(\mathbf{x})=\sum_{i=1}^{n}\mathbbm{1}(x_{i}>\epsilon)$, where $\epsilon$ is a small number. Then the loss is defined as $\ell_{y}(\mathbf{x})=|y-\operatorname{ND}(\mathbf{x})|$, which is non-differentiable.
@@ 4.2 Guidance via Stochastic Control
  Require: Loss function $\ell_{y}$, rule target $\mathbf{y}$, number of samples $n$.
@@ 5 Latent Diffusion Architecture
Model architecture. We first use a VAE model to encode short segments of piano rolls of shape $3\times 128\times 128$ into a latent space. Then we concatenate the latent codes and train a diffusion model to capture their joint distribution (Figure 2). For the VAE, we use the U-Net backbone following (Rombach et al., 2022). The training involves a denoising objective in conjunction with KL regularization: we introduce musically semantic perturbations (such as adding adjacent notes) to the data and train the model to revert to the original, unperturbed data. Both KL regularization and the denoising objective have proven indispensable for developing diffusion models with robust generative capabilities in subsequent stages.
@@ 6.1 Experimental Settings
Data. We train our model on several piano midi datasets that cover both classical and pop genres. The MAESTRO dataset (Hawthorne et al., 2019) has about 1200 pieces of classical piano performances with expressive dynamics, resulting in about 200 hours of music performance. In addition, we crawled about 14k MIDI files from Muscore in classical, religion, and soundtrack genres across all skill levels, yielding about 700 hours of data. We also used two Pop piano datasets: Pop1k7 (Hsiao et al., 2021) and Pop909 (Wang et al., 2020) that contain 108 hours and 60 hours of pop piano midi translated from audio respectively.
@@ 6.2 Unconditional Generation
Results. The evaluation results are in Table 2, highlighting the highest values (excluding GT) in bold. Our method achieves the highest average OA on all the datasets. The baselines are trained on individual dataset, and do not generalize well across datasets. MusicTr has the second-best overall rating for classical music (Maestro and Muscore), while it holds the lowest rating for pop music. CPW, on the other hand, ranks second in pop music but has the lowest rating in classical music. In contrast, our model delivers strong performance consistently across all the datasets.
Dataset GT MusicTr Remi CPW PolyDiff Ours
@@ 6.3 Individual Rule Guidance
Results. Table 3 shows the loss between the generated attributes and the target attributes, which measures the adherence to the rule. We make three main observations. First, our method significantly outperforms the other methods on non-differentiable rules (note density and chord progression). It achieves the lowest loss, without need for training any surrogate model, which is mandatory for classifier guidance and DPS-NN.
No Guidance $0.018\pm 0.010$ $2.486\pm 3.530$ $0.831\pm 0.142$
Classifier $0.005\pm 0.004$ $0.698\pm 0.587$ $0.723\pm 0.200$
DPS - NN $\mathbf{0.001\pm 0.002}$ $1.261\pm 2.340$ $0.414\pm 0.256$
DPS - Rule $0.010\pm 0.008$ $2.508\pm 2.798$
SCG (ours) $0.003\pm 0.004$ $\mathbf{0.131\pm 0.325}$ $\mathbf{0.273\pm 0.164}$
Table 3: Loss between the target and the generated attributes for individual rule guidance. SCG significantly improves the controllability of non-differentiable rules.
@@ 6.4 Composite Rule Guidance
No Guidance $0.018\pm 0.010$ $2.486\pm 3.530$ $0.831\pm 0.142$
Classifier $0.006\pm 0.006$ $0.822\pm 0.844$ $0.724\pm 0.205$
DPS-NN $0.004\pm 0.006$ $1.366\pm 2.265$ $0.661\pm 0.257$
SCG $0.014\pm 0.009$ $0.466\pm 0.648$ $0.446\pm 0.205$
SCG + DPS-NN $\mathbf{0.002\pm 0.007}$ $0.238\pm 0.531$ $0.313\pm 0.231$
SCG + Classifier $0.003\pm 0.005$ $\mathbf{0.148\pm 0.203}$ $\mathbf{0.284\pm 0.197}$
Table 4: Loss between the target and the generated attributes for individual rule guidance. SCG + Classifier achieves significantly lower losses for all three rules simultaneously.
We apply our method to generate samples that follow composite rules, following the same setup in Section 6.3, and assuming that the rule labels are conditionally independent given the sample. For classifier and DPS-NN, we train a surrogate model for each rule, and combine the gradient of each classifier to obtain the guidance term: $\Sigma_{i}\omega_{i}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{i}|\mathbf{x}_{t})$. For our method, we use a weighted loss function $\Sigma_{i}\omega_{i}\ell_{\mathbf{y}_{i}}(\mathbf{x})$ to select the best direction for each step. We set the weights on pitch histogram, note density and chord progression to be 40, 1, 1 respectively so that their loss is on the same order of magnitude.
The results are presented in Table 4. Our method achieves much lower loss on non-differentiable rules compared to other methods, similar to the case of individual rule guidance. However, it compromises control over the pitch histogram. Additionally, the loss associated with each rule is higher than in scenarios of individual rule guidance, which aligns with expectations. This increase in loss occurs because it is more challenging to identify a direction that satisfies multiple rules simultaneously, as opposed to a single rule, within the same computational budget.
@@ 6.5 Ablation Studies
Controllability and Quality Trade-Off. We observe a trade-off between controllability (measured by loss) and quality (measured by OA) in Table 5, where we guide the model to generate music following given note density. This is because there are more constraints on the generated music. For instance, a generative model could generate music that follows the note density exactly, but completely ignore the pitch of the notes. SCG achieves significant lower loss than the other guidance methods while maintaining reasonably high music quality. In addition, one can tune the balance between controllability and quality by tuning the number of samples at each step.
Method $n$ Loss $\downarrow$ OA $\uparrow$ Time (s)
Classifier 1 $0.698\pm 0.587$ $0.914\pm 0.006$ 47.8
DPS-NN 1 $1.261\pm 2.340$ $0.735\pm 0.012$ 109.3
4 $0.318\pm 0.770$ $\mathbf{0.895\pm 0.006}$ 277.7
SCG 8 $0.214\pm 0.368$ $0.877\pm 0.006$ 531.6
16 $\mathbf{0.131\pm 0.325}$ $0.880\pm 0.003$ 1242.6
4 $0.151\pm 0.298$ $\mathbf{0.906\pm 0.006}$ 301.9
Classifier + SCG 8 $0.098\pm 0.179$ $0.893\pm 0.004$ 555.6
16 $\mathbf{0.064\pm 0.159}$ $0.899\pm 0.007$ 1253.9
Table 5: Trade-offs between controllability, quality and computational time. $n$ refers to number of samples at each step.
@@
${}^{\dagger}$-800 750-200 $\mathbf{0.183\pm 0.341}$ $\mathbf{0.864\pm 0.005}$ 912.6
${}^{\dagger}$-700 750-300 $1.950\pm 1.344$ $0.737\pm 0.011$ 747.3
@@
Table 6: Impact of sampling strategy. The numbers that follow the method names are the total sampling steps. ${}^{\dagger}$: early stopping.
@@ 6.7 Examples of Our System as a Compositional Tool
To demonstrate how our system can be used effectively as a compositional tool, we provide links to three example videos, available through this website. For each video, a musician first indicated desired musical characteristics in terms of the rules (e.g. fairly sparse excerpt, following a simple I-V chord progression in C major, etc). The musician’s plan was to then loop this and use that as an accompaniment over which they would then improvise. The system generated 3-5 options (i.e. samples) for each such request, and the musician chose their preferred sample with which to work, from each set. They sent the generated MIDI output to a Disklavier piano and then recorded a second track over top of that. In the accompanying videos (linked below), you can see what the musician is playing, and everything “else” that you hear is from the material that was generated by the system.

### v4

@@ 1 Introduction
We are interested in developing methods for controllable symbolic music generation. There has been rapid progress in the development of modern generative models for symbolic music (Huang et al., 2018; Huang & Yang, 2020; Hsiao et al., 2021; Min et al., 2023). To facilitate interaction between human composers and these models, it is crucial for these models to adhere to specific musical rules, such as chord progression, during the composition process.
@@ 2 Related Works
The conceptualization of stochastic optimal control in diffusion models has spurred theoretical advancements and practical applications. Path integral theory (Kappen, 2005) provides an efficient way of solving stochastic optimal control problems. Zhang & Chen (2021) employed it to transform a simple Ornstein–Uhlenbeck process to a novel process whose target distribution matches given marginal distribution. Further extending this framework,  Berner et al. (2022); Vargas et al. (2023) established a novel link between stochastic optimal control problems and generative models, interconnected through stochastic differential equations.
@@ 3 Background
Classifier-free guidance (Ho & Salimans, 2022) avoids training classifiers by jointly training conditional and unconditional diffusion models, and combining their score estimates during sampling. The mixed score function becomes $(1+\omega)\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t}|\mathbf{y})-\omega\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$, where $\omega$ is the guidance strength. Despite easy implementation, it is expensive to extend classifier-free guidance to unknown or composite labels, because it requires re-training the diffusion model.
@@ 4.1 Rule Guidance Problem
A central challenge that we tackle is that many musical rules are non-differentiable, which makes sampling from Eq. 4 difficult. For instance, let $\mathbf{x}=[x_{1},x_{2},...,x_{n}]\in[0,1]^{n}$ be a vector where each $x_{i}$ represents the volume of a note, so that the note density is computed as $\operatorname{ND}(\mathbf{x})=\sum_{i=1}^{n}\mathbbm{1}(x_{i}>\epsilon)$, where $\epsilon$ is a small number. Let $\mathbf{y}=y$ be a scalar that represents the target note density. Then the loss is defined as $\ell_{\mathbf{y}}(\mathbf{x})=|y-\operatorname{ND}(\mathbf{x})|$, which is non-differentiable.
@@ 4.2 Guidance via Stochastic Control
  Require: Loss function $\ell_{y}$, rule target $\mathbf{y}$, number of samples $n$, forward process variances $\beta_{t}$, $\alpha_{t}:=1-\beta_{t}$, $\bar{\alpha}_{t}:=\prod_{s=1}^{t}\alpha_{s}$.
@@ 5 Latent Diffusion Architecture
Model architecture. We first use a VAE model to encode short segments of piano rolls of shape $3\times 128\times 128$ into a latent space. Then we concatenate the latent codes and train a diffusion model to capture their joint distribution (Figure 2). For the VAE, we use the same architecture following (Rombach et al., 2022). The training involves a denoising objective in conjunction with KL regularization: we introduce musically semantic perturbations (such as adding adjacent notes) to the data and train the model to revert to the original, unperturbed data. Both KL regularization and the denoising objective have proven indispensable for developing diffusion models with robust generative capabilities in subsequent stages.
@@ 6.1 Experimental Settings
Data. We train our model on several piano midi datasets that cover both classical and pop genres. The MAESTRO dataset (Hawthorne et al., 2019) has about 1200 pieces of classical piano performances with expressive dynamics, resulting in about 200 hours of music performance. In addition, we crawled about 14k MIDI files from Muscore in classical, religion, and soundtrack genres across all skill levels, yielding about 700 hours of data. We also used two Pop piano datasets: Pop1k7 (Hsiao et al., 2021) and Pop909 (Wang et al., 2020) that contain 108 hours and 60 hours of pop piano midi files.
@@ 6.2 Unconditional Generation
Results. The evaluation results are in Table 2, highlighting the highest values (excluding GT) in bold. Our method achieves the highest average OA on all the datasets. The detailed results for all 7 OA metrics are in Table 7, Appendix C. The baselines are trained on individual dataset, and do not generalize well across datasets. MusicTr has the second-best overall rating for classical music (Maestro and Muscore), while it holds the lowest rating for pop music. CPW, on the other hand, ranks second in pop music but has the lowest rating in classical music. In contrast, our model delivers strong performance consistently across all the datasets.
Dataset GT MusicTr Remi CPW PolyDiff SCG (ours)
@@
Method Loss $\downarrow$ (PH) OA $\uparrow$ (PH) Loss $\downarrow$ (ND) OA $\uparrow$ (ND) Loss $\downarrow$ (CP) OA $\uparrow$ (CP)
No Guidance $0.018\pm 0.010$ $0.842\pm 0.012$ $2.486\pm 3.530$ $0.830\pm 0.016$ $0.831\pm 0.142$ $\mathbf{0.854\pm 0.026}$
Classifier $0.005\pm 0.004$ $\underline{0.855\pm 0.020}$ $\underline{0.698\pm 0.587}$ $\mathbf{0.861\pm 0.025}$ $0.723\pm 0.200$ $0.850\pm 0.033$
DPS - NN $\mathbf{0.001\pm 0.002}$ $0.849\pm 0.018$ $1.261\pm 2.340$ $0.667\pm 0.113$ $\underline{0.414\pm 0.256}$ $0.839\pm 0.039$
DPS - Rule $0.010\pm 0.008$ $0.635\pm 0.006$ $2.508\pm 2.798$ $0.800\pm 0.080$
SCG (ours) $\underline{0.003\pm 0.004}$ $\mathbf{0.867\pm 0.005}$ $\mathbf{0.131\pm 0.325}$ $\underline{0.842\pm 0.031}$ $\mathbf{0.273\pm 0.1637}$ $\underline{0.851\pm 0.011}$
Table 3: Objective evaluation for individual rule guidance. Bottom 2 losses and top 2 OA metrics are highlighted. SCG significantly improves the controllability of non-differentiable rules.
@@ 6.3 Individual Rule Guidance
Metrics. We evaluate conditional generation performance by two metrics: loss and OA. Loss reflects controllability: whether the generated samples follow the target rules. For pitch histogram and note density, we use L2 loss, and for chord progression, we use 0-1 loss. However, a low loss does not necessarily indicates good quality music. For instance, a music piece that follows the note density requirement may have random pitch and does not sound good. Therefore, we use the OA to measure how close the generated music distribution and the ground truth music distribution (with similar target attributes as the generated music) when projected to some music features space. Please see Appendix D.3 for the detailed evaluation setup.
Results. Table 3 shows the results. We make three main observations. First, our method significantly outperforms the other methods in generating samples that follow the non-differentiable rules (note density and chord progression). It achieves the lowest loss, without need for training any surrogate model, which is mandatory for classifier guidance and DPS-NN. However, it sacrifices OA a bit. This is because over-optimizing for one attribute will overlook the other attributes. Reducing the number of samples used for SCG can lead to better OA at the cost of a higher loss (See Section 6.5 on this trade-off).
@@ 6.4 Composite Rule Guidance
We apply our method to generate samples that follow composite rules, following the same setup in Section 6.3, and assuming that the rule labels are conditionally independent given the sample. For classifier and DPS-NN, we train a surrogate model for each rule, and combine the gradient of each classifier to obtain the guidance term: $\Sigma_{i}\omega_{i}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{y}_{i}|\mathbf{x}_{t})$. For our method, we use a weighted loss function $\Sigma_{i}\omega_{i}\ell_{\mathbf{y}_{i}}(\mathbf{x})$ to select the best direction for each step. We set the weights on pitch histogram, note density and chord progression to be 40, 1, 1 respectively so that their loss is on the same order of magnitude. In addition, we compared with four established conditional music generation baselines: Retrieval, Figaro-expert, Figaro-expert-learned (von Rütte et al., 2022) and MuseCOCO (Lu et al., 2023). The Retrieval baseline is: given a set of target attributes, we find the sample within the datasets that has the closest attributes to the target attributes. This serves as an oracle baseline.
Method PH $\downarrow$ ND $\downarrow$ CP $\downarrow$ OA $\uparrow$
Retrieval $0.006\pm 0.005$ $0.433\pm 1.068$ $0.556\pm 0.182$ $\mathbf{0.886\pm 0.005}$
Figaro expert $0.007\pm 0.007$ $2.303\pm 2.256$ $0.761\pm 0.187$ $0.857\pm 0.047$
Figaro expert+learned $0.006\pm 0.009$ $1.489\pm 2.737$ $0.726\pm 0.214$ $0.883\pm 0.008$
MuseCoCo $0.040\pm 0.026$ $2.734\pm 3.551$ $0.821\pm 0.163$ $0.753\pm 0.038$
No Guidance $0.018\pm 0.010$ $2.486\pm 3.530$ $0.831\pm 0.142$ $0.803\pm 0.096$
Classifier $0.006\pm 0.006$ $0.822\pm 0.844$ $0.724\pm 0.205$ $0.859\pm 0.026$
DPS-NN $0.004\pm 0.006$ $1.366\pm 2.265$ $0.661\pm 0.257$ $0.752\pm 0.079$
SCG $0.014\pm 0.009$ $0.466\pm 0.648$ $0.446\pm 0.205$ $\mathbf{0.874\pm 0.007}$
DPS-NN + SCG $\mathbf{0.002\pm 0.007}$ $0.238\pm 0.531$ $0.313\pm 0.231$ $0.781\pm 0.084$
Classifier + SCG $0.003\pm 0.005$ $\mathbf{0.148\pm 0.203}$ $\mathbf{0.284\pm 0.197}$ $0.864\pm 0.010$
Table 4: Objective evaluation for composite rule guidance. SCG + Classifier achieves significantly lower losses for all three rules simultaneously with a high OA score.
The results are presented in Table 4. We can see that, overall, our method achieves the lowest loss among all the baselines. The OA ranking is also higher than that of individual rule guidance, because guiding the generation with three attributes prevents over-optimizing for one attribute. There are two baselines that have higher OA than us: retrieval and Figaro-expert-learned. It is as expected that retrieval has the highest OA because the samples are selected from the dataset rather than being generated. Figaro-expert-learned also achieves high OA, because it is a style transfer model that needs to take in source music rather than a generative model. The generated music is conditioned on the latent representation of the source music. Therefore, it is not a fair comparison but we still put it here for reference. MuseCOCO cannot generate well following the target rules because it does not support fine-grained control.
Among the diffusion guidance methods, our method achieves much lower loss on non-differentiable rules compared to other methods, similar to the case of individual rule guidance. However, it compromises control over the pitch histogram. Additionally, the loss associated with each rule is higher than in scenarios of individual rule guidance, which aligns with expectations. This increase in loss occurs because it is more challenging to identify a direction that satisfies multiple rules simultaneously, as opposed to a single rule, within the same computational budget.
@@ 6.5 Ablation Studies
Controllability and OA Trade-Off. We observe a trade-off between controllability (measured by loss) and OA (used as a necessary, albeit not sufficient, indicator of music quality.) in Table 5, where we guide the model to generate music following given note density. To better evaluate the trade-off, we compute 2 OA metrics, one uses the full dataset as reference (denoted by ‘OA full’), the other uses the data that comply with the desired rule (denoted by ‘OA’ as in previous sections). The motivation of ‘OA full’ stems from our approach of extracting note density from a diverse selection of sources within the dataset for conditional generation. If the generated pieces precisely mirror their source, they can achieve a low loss because they comply with the rule perfectly, and have a high OA because their sources are taken from the dataset (the ‘source’ row in Table 5).
We highlight the highest OA for each group of methods (the baselines, SCG and classifier+SCG with different number of samples). Note that the main difference between these two metrics is for ‘No Guidance’, where ‘OA cluster’ is much worse than ‘OA full’. This is because the ‘OA cluster’ metric rewards controllability more. However the controllability and OA trade-off for SCG related methods are consistent among the two OA metrics. We think that this trade-off is caused by over-optimizing over some constraints. For instance, a generative model could generate music that follows the note density exactly, but completely ignore the pitch of the notes. One can balance controllability and OA by tuning the number of samples at each step.
Method $n$ Loss $\downarrow$ OA full $\uparrow$ OA $\uparrow$ Time (s)
No Guidance - $2.486\pm 3.530$ $0.918\pm 0.005$ $0.830\pm 0.016$ 25.4
Source - 0 $\mathbf{0.923\pm 0.008}$
Classifier 1 $0.698\pm 0.587$ $0.914\pm 0.006$ $\mathbf{0.861\pm 0.025}$ 47.8
DPS-NN 1 $1.261\pm 2.340$ $0.735\pm 0.012$ $0.667\pm 0.113$ 109.3
4 $0.318\pm 0.770$ $\mathbf{0.895\pm 0.006}$ $\mathbf{0.873\pm 0.023}$ 277.7
SCG 8 $0.214\pm 0.368$ $0.877\pm 0.006$ $0.847\pm 0.014$ 531.6
16 $\mathbf{0.131\pm 0.325}$ $0.880\pm 0.003$ $0.842\pm 0.031$ 1242.6
4 $0.151\pm 0.298$ $\mathbf{0.906\pm 0.006}$ $\mathbf{0.861\pm 0.011}$ 301.9
Classifier + SCG 8 $0.098\pm 0.179$ $0.893\pm 0.004$ $0.839\pm 0.024$ 555.6
16 $\mathbf{0.064\pm 0.159}$ $0.899\pm 0.007$ $0.849\pm 0.018$ 1253.9
Table 5: Trade-offs between controllability, OA and computational time. $n$ refers to number of samples at each step.
@@
DDPM-800 750-200 $\mathbf{0.183\pm 0.341}$ $\mathbf{0.864\pm 0.005}$ 912.6
DDPM-700 750-300 $1.950\pm 1.344$ $0.737\pm 0.011$ 747.3
@@
Table 6: Impact of sampling strategy. The numbers that follow the method names are the total sampling steps. : early stopping.
@@ 6.7 Examples of Our System as a Compositional Tool
To demonstrate how our system can be used effectively as a compositional tool, we provide links to three example videos, available through this website. For each video, a musician first indicated desired musical characteristics in terms of the rules (e.g. fairly sparse excerpt, following a simple I-V chord progression in C major, etc). The musician’s plan was to then loop this and use that as an accompaniment over which they would then improvise. The system generated 3-5 options (i.e. samples) for each such request, and the musician chose their preferred sample with which to work, from each set. They sent the generated MIDI output to a Disklavier piano and then recorded a second track over top of that. In the accompanying videos, the musician is playing along with the generated music by our approach.
@@ 7 Conclusion and Future directions
In principle, the SCG algorithm introduced in this paper extends beyond the realm of symbolic music generation. Its capability to enforce diffusion models to follow non-differentiable constraints makes it suitable for diverse fields, as long as the constraints can be programmed and one can define a loss on how well the constraints are satisfied. For instances, in protein design, one can write a function to return how specific topological constraints are satisfied. In astronomy imaging, one can use black-box physics simulators as the rule function. We believe it is an exciting future direction to extend this algorithm to a broader scope.
@@ Acknowledgements
Impact Statement
This paper presents work whose goal is to advance the field of generative modeling for symbolic music. While there are numerous potential societal implications associated with our work, we believe none require specific emphasis in this context.
@@ C.1 Unconditional Generation
In addition, we check if the models are copying from the dataset. To do so, we randomly pick 50 generated samples for each method and 2000 samples from the training dataset, and identify the closest musical piece to a generated MIDI file from the dataset. Specifically, we extract and compare key features from each file, including pitch, velocity, duration of notes, and overall note density. We report the average matching score of these four features in Table 8. As we can see, despite our method achieving the highest OA, it has the second lowest matching score, indicating that the OA improvement is not from copying from the dataset.
Method Matching Score OA
MusicTr $0.6273\pm 0.0336$ $0.903\pm 0.005$
Remi $0.6386\pm 0.0210$ $0.866\pm 0.004$
CPW $0.6439\pm 0.0134$ $0.899\pm 0.005$
PolyDiff $0.6372\pm 0.0097$ $0.883\pm 0.004$
Ours $0.6337\pm 0.0185$ $0.943\pm 0.003$
Table 8: Test if the model is copying samples from the dataset by evaluating Matching Score and OA.